is 43 a prime number?
Found 2 solutions by htmentor, ewatrrr:
Answer by htmentor(1343)

We need to check if the number is divisible by anything other than 1 and 43.
Since it is odd, it is not divisible by 2, which means it is not divisible by 4, 6, or 8 (and if it were, it wouldn't be prime, anyway!)
Is it divisible by 3? No. 43/3 = 14, remainder 1
It is not divisible by 5, since it does not end in 0 or 5.
Is it divisible by 7? No. 43/7 = 6, remainder 1
Since it is not divisible by 3, it is not divisible by 9.
Therefore, the number is prime.
Note that we could have stopped checking at 7, since 7*7 = 49 > 43.

Answer by ewatrrr(24785)


Hi
Prime numbers between 1 and 100 are as follows:
2 3 5 7 11 13 17 19 23 29 31 37 41 43 47 53 59 61 67 71 73 79 83 89 97
is 43 a prime number? YES, Note: not divisible by any number other than itself.